40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 7183     Entry time: Tue Aug 14 21:01:51 2012
Author: rana 
Type: Update 
Category: PEM 
Subject: BLRMS 

Screenshot-Untitled_Window.png

I fixed up the seismic.stp file for the StripTool display:

  1. All BLRMS channels now have a y-axis range of 3 decades. So they all are displaying the same relative changes.
  2. So the 0.01-0.1 Hz band which is all over the place is real, sort of. Masha says that it is due to the seismometer signal being dominated by noise below 0.1 Hz. She is going to fix this somehow.
  3. I changed the samping time from 1 sec. to 10 sec. to make the traces less fuzzy.
  4. We (Masha / Liz) should harmonize the colors of this file with what's on the summary pages.
ELOG V3.1.3-